OpenStudy (anonymous):

Absolute value equation Can you help me? |-4m|=64 |7c+3|=c-1 |a-4|-20=-3a Thank you!

OpenStudy (anonymous):

Can somebody help?

OpenStudy (blurbendy):

|-4m| = 64 4m = 64 m = ??? then consider -4m = 64 So you'll end up with two answers.

OpenStudy (anonymous):

Your first question answer |-4m| = 64 m can be minus or plus doesnt matter so m=16 or m=-16 Second one |7c+3|=c-1 c has to bigger than 1 if not equation will be not true so 7c+3=c-1 when u look this equation c will be minus this is not possible so this question is not possible to solve |a-4|-20=-3a for a<4 4-a-20=-3a -16=-2a a=8 but a should be smallert than 4 so its not true for a>4 a-4-20=-3a 4a=24 a=6 yes this is the what we are looking for so the answer is 6